a system releases 622kJ of heat and does 105 kJ of word on the surroundings. what is the change in internal energy of the system?


1
Expert's answer
2021-11-23T10:35:01-0500

According to the first law of thermodynamics



"\Delta\>u=q+w"


"\Delta\>u=-622-105\\=-727kJ"